OpenStudy (clara1223):

solve for x: 4^(-x-3)+3=35 a) x=15/2 b) x=-8 c) x=-1/2 d) x=11/2 e) x=-11/2

OpenStudy (clara1223):

\[4^{-x-3}+3=35\]

OpenStudy (decentnabeel):

\[4^{\left(-x-3\right)}+3=35\quad :\quad x=-\frac{11}{2}\]

OpenStudy (decentnabeel):

answer is option e

OpenStudy (clara1223):

@DecentNabeel please explain

OpenStudy (decentnabeel):

ok 4^(-x-3)+3=35 first subtract 3 on both sides

OpenStudy (decentnabeel):

4^(-x-3)+3-3=35-3 4^(-x-3)=32

OpenStudy (decentnabeel):

\[\left(2^2\right)^{-x-3}=2^5\] \[\left(2^2\right)^{-x-3}=2^5\] \[\mathrm{Use\:the\:following\:exponent\:property}:\quad \left(a^n\right)^m=a^{n\cdot m}\] \[\left(2^2\right)^{-x-3}=2^{2\left(-x-3\right)}\] \[2^{2\left(-x-3\right)}=2^5\] \[\mathrm{If\:}a^{f\left(x\right)}=a^{g\left(x\right)}\mathrm{,\:then\:}f\left(x\right)=g\left(x\right)\] \[\mathrm{Solve\:}\:2\left(-x-3\right)=5:\quad x=-\frac{11}{2}\]

OpenStudy (decentnabeel):

are you understand @clara1223

OpenStudy (clara1223):

yes! thanks so much @DecentNabeel

OpenStudy (decentnabeel):

you are wellcome @clara1223